What percent of the total cost is represented by each category of expenses?
A contractor built a warehouse complex for the followoing cost: land $12,000, concete and steel $34,500, plumbing and electrical $48,990, general carpentry and roof $42,340, and other expenses $34,220

Answer by mananth(16949)

A contractor built a warehouse complex for the followoing cost: land $12,000, concete and steel $34,500, plumbing and electrical $48,990, general carpentry and roof $42,340, and other expenses $34,220
...
land ...........................$12,000,
concrete and steel ..............$34,500,
plumbing and electrical........ $48,990,
general carpentry and roof......$42,340,
and other expenses..............$34,220
---------------------------------------
TOTAL =.......................$172,050
..
Land 12,000/172,050 * 100= 6.97%
...
concrete and steel $34,500/172,050 *100=20.05%
...
plumbing and electrical $48,990/172,050*100= 28.47%
general carpentry and roof$42,340/172050*100= 24.6%
and other expenses$34,220/172050*100=19.9%